# CMM-Math: A Chinese Multimodal Math Dataset To Evaluate and Enhance the Mathematics Reasoning of Large Multimodal Models

### v2

@@ Abstract
Large language models (LLMs) have obtained promising results in mathematical reasoning, which is a foundational skill for human intelligence. Most previous studies focus on improving and measuring the performance of LLMs based on textual math reasoning datasets (e.g., MATH, GSM8K). Recently, a few researchers have released English multimodal math datasets (e.g., MATHVISTA and MATH-V) to evaluate the effectiveness of large multimodal models (LMMs). In this paper, we release a Chinese multimodal math (CMM-Math) dataset, including benchmark and training parts, to evaluate and enhance the mathematical reasoning of LMMs. CMM-Math contains over 28,000 high-quality samples, featuring a variety of problem types (e.g., multiple-choice, fill-in-the-blank, and so on) with detailed solutions across 12 grade levels from elementary to high school in China. Specifically, the visual context may be present in the questions or opinions, which makes this dataset more challenging. Through comprehensive analysis, we discover that state-of-the-art LMMs on the CMM-Math dataset face challenges, emphasizing the necessity for further improvements in LMM development. We also propose a Multimodal Mathematical LMM (Math-LMM) to handle the problems with mixed input of multiple images and text segments. We train our model using three stages, including foundational pre-training, foundational fine-tuning, and mathematical fine-tuning. The extensive experiments indicate that our model effectively improves math reasoning performance by comparing it with the SOTA LMMs over three multimodal mathematical datasets.
Figure 1: Some examples in our CMM-Math datasets.
Introduction
@@
The recent advent of large language models (LLMs), like OpenAI’s GPT-4 (OpenAI 2023) and LLaMA (Touvron et al. 2023a, b), have demonstrated remarkable success across math problem solving, which is an important skill for human intelligence (Liu et al. 2023b). Several studies utilize chain-of-thought (CoT) (Wei et al. 2022) to solve the problems step by step (Chen et al. 2022; Lu et al. 2023; Zhang et al. 2023; Fu et al. 2023). Also, tools and programs are integrated with LLMs to reduce the hallucination problem of math reasoning (Gao et al. 2023a; Drori et al. 2022; He-Yueya et al. 2023). LLEMMA (Azerbayev et al. 2023) is a math-specific LLM that is fine-tuned on large-scale mathematical datasets, which obtains great abilities in generating textual solutions for mathematical problems and formulating formal proofs.
@@
Furthermore, to evaluate and improve the mathematical performance of LLMs, some textual math datasets are released (Mishra et al. 2022; Frieder et al. 2024). For instance, Hendrycks et al. proposed a MATH dataset that contains 12,500 challenging competition mathematics problems with step-by-step solutions. GSM8K (Cobbe et al. 2021) is a dataset of 8.5K high-quality linguistically diverse grade school math word problems. However, these studies mainly focus on textual information while the visual contexts of the problems are ignored by the previous literature.
@@
Recently, various large multimodal models (LMMs) like GPT-4V (OpenAI 2023), LLaVA (Liu et al. 2024a), CogVLM (Wang et al. 2023) and Gemini (Team et al. 2024) are proposed for improving the performance of multimodal reasoning. To measure the mathematic reasoning abilities of LLMs, some researchers released multimodal math benchmarks. MMMU (Yue et al. 2024) is a popular multimodal dataset that mainly focuses on visual recognition with few problems about simple mathematical reasoning. Additionally, the multimodal datasets designed for mathematical reasoning are also proposed (Lu et al. 2024; Wang et al. 2024). Lu et al. (2024) proposed the first math-specific English benchmark dataset, MATHVISTA, to measure the performance of the LMMs. This dataset includes seven mathematical reasoning types with a diverse array of visual contexts across five primary tasks. Additionally, Wang et al. (2024) organized a math evaluation benchmark, MATH-V, which contains 3,040 mathematical problems across 16 knowledge concepts.
@@
Unlike these studies, the unique characteristics of our Chinese Mathematical (CMM-Math) dataset are listed as follows. First, the existing multimodal mathematical datasets are in English, while our CMM-Math focuses on Chinese. Second, we release a benchmark dataset to measure performance and contribute a training dataset to fine-tune LMMs, facilitating future research and increasing accuracy in multimodal mathematical reasoning. Third, our dataset is more complex, with problems that may contain multiple figures in questions or options, as shown in the first example in Figure 1. Moreover, it contains more question types, including single-choice questions, multiple-choice questions, fill-in-the-blank questions, and essay questions. Fourth, our large-scale CMM-Math datasets contain more than 28,000 samples across 12 grade levels from elementary to high school. To our knowledge, is currently the largest Chinese mathematical dataset available.
@@
We evaluate the state-of-the-art open-source and close-source LMMs on our CMM-Math dataset. The results show that it is a challenge for the current foundation LMMs to handle multimodal mathematical reasoning tasks. Furthermore, we propose a math-specific LMM (Math-LMM) with a mixed instruction to handle the problems with multiple images. We train Math-LMM using three phases: foundational pre-training that aligns visual information with LLM, foundational fine-tuning that captures the abilities of task solving, and mathematical fine-tuning that learns mathematical reasoning.
@@
  • We collect and release a high-quality multimodal mathematical dataset, CMM-Math, which consists of evaluation and training datasets to measure and improve the performance of the current LMMs.
@@
  • We conduct a comprehensive evaluation of our Math-LMM model and the existing LMMs on our CMM-Math dataset. The results show that the performance of the existing SOTA methods is limited. Our model performs well over MATHVISTA and Math-V datasets with only about 8B parameters.
Figure 2: The performance of LMMs across 13 subjects
Figure 3: The performance of LMMs across 12 levels
Related Works
Mathematical Dataset
@@
Various datasets (Mishra et al. 2022; Cobbe et al. 2021; Hendrycks et al. 2020, 2021a; Frieder et al. 2024)) have been introduced to measure the mathematical capabilities of large language models (LLMs). For instance, GSM8K (Cobbe et al. 2021) and MATH (Hendrycks et al. 2021a) are two popular textual datasets to comprehensively evaluate the LLMs. These datasets (Liu et al. 2023b) can test the language processing abilities in areas such as mathematical reasoning, computation, and theorem proving, but they do not effectively assess the ability of large multimodal models to solve math problems that require visual modalities, such as geometry problems, properties of function graphs, etc.
@@
In the past two years, specialized multimodal datasets for the mathematics domain have emerged to assess the mathematical capabilities of large multimodal models, such as MMMU (Yue et al. 2024), MATHVISTA (Lu et al. 2024), and MATH-V (Wang et al. 2024). MMMU (Yue et al. 2024) primarily evaluates the model’s visual recognition abilities, with only a few questions involving simple mathematical reasoning. MATHVISTA (Lu et al. 2024) consolidates and transforms existing FQA (Figure Question Answering), GPS (Geometry Problem Solving), MWP (Math Word Problems), TQA (Textbook Question Answering), and VQA(Visual Question Answering) datasets, enabling tests of basic mathematical reasoning, but it covers fewer math concepts and has a limited variety of question types, many of which can be solved with VQA capabilities alone. MATH-V (Wang et al. 2024) compiles questions from math competitions like Math Kangaroo, UK contests at various levels (Grey, Pink, Junior, Senior), and US competitions (AMC 8, 10, 12) and invitational events (AIME). This dataset evaluates models across multiple educational levels and topics.
@@
Unlike these datasets, our CMM-Math dataset not only offers evaluations covering multiple grades and topics but also includes both training and evaluation datasets. Furthermore, the problem in our dataset may contain multiple images with detailed solution explanations across a variety of problem types. We have also provided a large amount of data that contains only the text modality, which, to our knowledge, is currently the largest Chinese mathematical dataset available.
Large Multimodal Model
@@
With the introduction of GPT-4V (OpenAI 2023), an increasing number of versatile multimodal large models have been proposed. Commonly, these models aim to handle both text and image tasks simultaneously. They employ a tokenizer and a visual encoder (such as CLIP (Radford et al. 2021)) to encode textual and visual information separately, then concatenate the encoded vectors into a unified input that is fed into the large language model. Notable large multimodal models include LLaVa (Liu et al. 2024a), Qwen-VL (Bai et al. 2023b), Gemini (Team et al. 2024), InternLMX Composer-VL (Dong et al. 2024; Zhang et al. 2024a), and GPT-4o. The CMM-Math can provide detailed and specialized comparisons of these models’ multimodal mathematical capabilities. Furthermore, we propose a math-specific LMM for multimodal mathematical reasoning, which obtains new state-of-the-art performance.
Table 1: Key statistics of CMM-Math.
@@
Dataset
Dataset Construction
For a lack of Chinese multimodal mathematical datasets, we collect and release CMM-Math dataset. Our benchmark data collection is divided into three stages. In the first stage, we collect more than 10,000 exam papers with a total of 12 grades of mathematics test problems from primary school to high school in China. Each test paper includes multiple-choice, fill-in-the-blank, analysis, and other problems, with problems containing visual input or plain text. Most of these papers provide the solution, grade, Since the test problems are in PDF format, we use the Mathpix API [^1] https://mathpix.com to extract text and images in a markdown format and download the extracted images locally. In the second stage, we phrase the problems from the markdown text using a JSON format, including fields such as question type, modality, question, option, answer, solution, etc. In order to improve the quality of the data, we conducted multiple rounds of validation. Firstly, we will examine the issue of image recognition. Correct text that Mathpix fails to recognize accurately (recognizing text or mathematical formulas as images) and images that are incorrectly recognized. Secondly, detect potential issues with each field, including mismatched question types and lengthy parsing. In the third stage, we divided the problem into 13 different topics. Due to the large amount of data and the high cost of manual annotation, we chose three large models, GPT-4o, Gemini, and Qwen-vl-max to vote on the topics and assign them to the topics based on the principle of minority obeying majority. Finally, we ask students to check the problems with the format and context.
Table 2: Detailed statistics of CMM-Math datasets.
@@
Multimodal Training Evaluation Language #Number Type
MATH (Hendrycks et al. 2021a) EN 12,500 Fill-in-the-blank questions
GSM8K (Cobbe et al. 2021) EN 8,500 Fill-in-the-blank questions
MathQA (Amini et al. 2019) EN 37,259 Fill-in-the-blank questions
MMMU (Yue et al. 2024) EN 11,500 Choice and fill-in-the-blank questions
MATHVISTA (Lu et al. 2024) EN 6,141 Choice and fill-in-the-blank questions
MATH-V (Wang et al. 2024) EN 3,040 Choice and fill-in-the-blank questions
CMM-Math (Ours) CN 28,249 Choice, fill-in-the-blank, analysis and yes-no questions
Table 3: Comparison with existing datasets. EN and CN mean English and Chinese.
Figure 4: The diagram of our Math-LMM framework.
Dataset Analysis
Our CMM-Math dataset contains 28,069 problems with rich textual and visual information (See Table 1 and Table 2). Among them, there are 21,200 textual problems and 6,869 multimodal problems, which can be divided into choice, fill-in-the-blank, yes-no, and analysis problems. We also split the dataset into 12 levels, corresponding to the basic education stage from first grade of primary school to third grade of high school, to ensure the applicability and reference value of the dataset in educational practice. CMM-Math contains 13 knowledge points, covering most of the mathematical fields encountered in middle and high school, particularly logic, algebra, counting, arithmetic, combinatorics, graph theory, topology, statistics, solid geometry, metric geometry, analytic geometry, descriptive geometry, combinatorial geometry, and transformation geometry.
@@
Comparison with Existing Datasets
@@
To clarify the characteristics of our CMM-Math, we compare it with multiple existing mathematical benchmarks, including textual datasets (e.g., MATH, GSM8K, MathQA) and multimodal datasets (e.g., MMMU, MATHVISTA, MATH-V), as shown in Table 3. First, most of these datasets are in English while we focus on Chinese to measure LMMs more comprehensively. Second, our CMM-Math is a multimodal dataset for mathematical reasoning, where each problem may contain multiple images, designed for LMMs. Textual MATH and GSM8K datasets are used to evaluate the performance of LLMs. Third, our dataset contains the evaluation and training datasets. The existing multimodal datasets evaluate the power of LMMs while ignoring how to improve their performance. Fourth, our datasets are diverse, with rich subjects and problem types across 12 grades, which can more deeply distinguish models’ performance in different mathematical contexts. CMM-Math provides 11 knowledge points and multiple question types, including choice, yes-no, computation, fill-in-the-blank, and analysis. In addition, the CMM-Math problems are derived from real math tests in primary, middle, and high school, covering a wider range of grades than the GSM8K, which only includes primary school knowledge. CMM-Math can better test the problem-solving ability of the model at different stages. Finally, our dataset has detailed analysis and rich question stem content, which can enhance the mathematical reasoning ability of the model.
Our Proposed Method

### v3

@@ Abstract.
Large language models (LLMs) have obtained promising results in mathematical reasoning, which is a foundational skill for human intelligence. Most previous studies focus on improving or measuring the performance of LLMs based on textual math datasets (e.g., MATH, GSM8K). Recently, a few researchers have released English multimodal math datasets (e.g., MATHVISTA and MATH-V) to evaluate the effectiveness of large multimodal models (LMMs). In this paper, we release a Chinese multimodal math (CMM-Math) dataset, including benchmark and training parts, to evaluate and enhance the mathematical reasoning of LMMs. CMM-Math contains over 28,000 high-quality samples, featuring a variety of problem types (e.g., choice, fill-in-the-blank, analysis) with detailed solutions across 12 grade levels from elementary to high school in China. Each problem may contain multiple images, and the visual context may be present in the questions or opinions, which makes this dataset more challenging. Our comprehensive analysis reveals that state-of-the-art LMMs on the CMM-Math dataset face challenges, emphasizing the necessity for further improvements in LMM development. We also propose a Multimodal Mathematical LMM (Math-LMM) to handle the problems with mixed input of multiple images and text segments. The Math-LMM is trained using three stages: foundational pre-training, foundational fine-tuning, and mathematical fine-tuning. The extensive experiments indicate that our model effectively improves math reasoning performance by comparing it with the SOTA LMMs over three multimodal mathematical datasets. Our datasets, codes, and weights will be released on GitHub.
Mathematical Reasoning, Large Multimodal Models, Benchmark, Chinese
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; 28 April - 2 May, 2025; Sydney, Australiaisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems World Wide Webccs: Computing methodologies Natural language processingccs: Networks Network algorithms
Figure 1. The performance of large multimodal models (LMMs) across 12 difficulty levels in our CMM-Math, where Level 1 is the easiest and Level 12 is the most difficult.
[…]
The performance of LMMs across 12 levels.
Figure 2. Some examples in our CMM-Math datasets.
1. Introduction
@@
The recent advent of large language models (LLMs), like OpenAI’s GPT-4 (OpenAI, 2023) and LLaMA (Touvron et al., 2023), have demonstrated remarkable success across math problem solving, which is an important skill for human intelligence (Liu et al., 2023). Several studies utilize chain-of-thought (CoT) (Wei et al., 2022) to solve the problems step by step (Zhang et al., 2023). Also, tools and programs are integrated with LLMs to reduce the hallucination problem of math reasoning (Gao et al., 2023b). LLEMMA (Azerbayev et al., 2023) is a math-specific LLM that is fine-tuned on large-scale mathematical datasets, which obtains great abilities in generating textual solutions for mathematical problems and formulating formal proofs.
@@
Furthermore, to evaluate and improve the mathematical performance of LLMs, some textual math datasets are released (Mishra et al., 2022; Frieder et al., 2024). For instance, Hendrycks et al. ([n. d.]) proposed a MATH dataset that contains 12,500 challenging competition mathematics problems with step-by-step solutions. GSM8K (Cobbe et al., 2021) is a dataset of 8.5K high-quality linguistically diverse grade school math word problems. However, these studies mainly focus on textual information while the visual contexts of the problems are not well studied by the previous literature.
@@
Recently, various large multimodal models (LMMs) like GPT-4V (OpenAI, 2023), LLaVA (Liu et al., 2024b), CogVLM (Wang et al., 2023) and Gemini (Team et al., 2024) are proposed for improving the performance of multimodal reasoning. To measure the mathematic reasoning abilities of LMMs, some researchers released multimodal math benchmarks. MMMU (Yue et al., 2024) is a popular multimodal dataset that mainly focuses on visual recognition with few problems about simple mathematical reasoning. Additionally, the multimodal datasets designed for mathematical reasoning are also proposed (Lu et al., 2024; Wang et al., 2024b). Lu et al. (2024) proposed the first math-specific English benchmark dataset, MATHVISTA, to measure the performance of the LMMs. This dataset includes seven mathematical reasoning types with diverse visual contexts across five primary tasks. Then, Qiao et al. (2024) collected 6.5K visual math problems to explore the problem-solving principles beyond the end-to-end performance. Additionally, Wang et al. (2024b) organized a math evaluation benchmark, MATH-V, which contains 3,040 mathematical problems across 16 knowledge concepts. These studies mainly focus on measuring the performance of mathematical reasoning for the existing LMMs in English.
@@
Unlike these studies, the unique characteristics of our Chinese Mathematical (CMM-Math) dataset are listed as follows. First, the existing multimodal mathematical datasets are in English, while our CMM-Math focuses on Chinese. Second, we release a benchmark dataset to measure performance and contribute a training dataset to fine-tune LMMs, facilitating future research and increasing accuracy in multimodal mathematical reasoning. Third, our dataset is more complex, with problems that may contain multiple figures in questions or options, as shown in the first example in Figure 2. Moreover, it contains more question types, including choice problems, fill-in-the-blank problems, yes-no problems, and analysis problems. Fourth, our large-scale CMM-Math datasets contain more than 28,000 samples across 12 grade levels from elementary to high school, to ensure the completeness and diversity of the dataset.
@@
We evaluate the state-of-the-art open-source and close-source LMMs on our CMM-Math dataset. The results show that it is a challenge for the current foundation LMMs to handle multimodal mathematical reasoning tasks. Furthermore, we propose a math-specific LMM (Math-LMM) with a mixed instruction to handle the problems with multiple images. We train Math-LMM using three phases: foundational pre-training that aligns visual information with LLM, foundational fine-tuning that captures the abilities of task solving, and mathematical fine-tuning that learns mathematical reasoning. We evaluate the performance of Math-LMM over three benchmark datasets and the results show our model outperforms the open-source LMMs in most cases.
@@
  • We collect and release a high-quality Chinese multimodal mathematical dataset, CMM-Math, which consists of evaluation and training datasets to measure and improve the performance of the current LMMs.
@@
  • We conduct a comprehensive evaluation of the existing LMMs and our Math-LMM model on the CMM-Math dataset. The results show that multimodal mathematical reasoning is still a great challenge for LLMs. Our model outperforms the strong open-source LMMs over CMM-Math, MATHVISTA and Math-V datasets in most cases.
2. Related Works
2.1. Mathematical Dataset
@@
Various datasets (Mishra et al., 2022; Cobbe et al., 2021; Hendrycks et al., 2021a; Frieder et al., 2024) have been introduced to measure the mathematical capabilities of large language models (LLMs). For instance, GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021a) are two popular textual datasets to comprehensively evaluate the LLMs. These datasets (Liu et al., 2023) can test the language processing abilities in areas such as mathematical reasoning, computation, and theorem proving, but they do not effectively assess the ability of large multimodal models to solve math problems that require visual modalities, such as geometry problems, properties of function graphs, etc.
@@
In the past two years, specialized multimodal datasets for the mathematics domain have emerged to assess the mathematical capabilities of large multimodal models, such as MMMU (Yue et al., 2024), MATHVISTA (Lu et al., 2024), We-Math (Qiao et al., 2024) and MATH-V (Wang et al., 2024b). MMMU (Yue et al., 2024) primarily evaluates the model’s visual recognition abilities, with only a few questions involving simple mathematical reasoning. MATHVISTA (Lu et al., 2024) consolidates and transforms existing FQA (Figure Question Answering), GPS (Geometry Problem Solving), MWP (Math Word Problems), TQA (Textbook Question Answering), and VQA(Visual Question Answering) datasets, enabling tests of basic mathematical reasoning, but it covers fewer math concepts and has a limited variety of question types, many of which can be solved with VQA capabilities alone. MATH-V (Wang et al., 2024b) compiles questions from math competitions like Math Kangaroo, UK contests at various levels (Grey, Pink, Junior, Senior), and US competitions (AMC 8, 10, 12) and invitational events (AIME). This dataset evaluates models across multiple educational levels and topics.
@@
Unlike these datasets, our CMM-Math dataset not only offers evaluations covering multiple grades and topics but also includes both training and evaluation datasets. Furthermore, the problem in our dataset may contain multiple images with detailed solution explanations across a variety of problem types. We have also provided a large amount of Chinese multimodal problems with multiple images across 12 grade levels from elementary to high school.
2.2. Large Multimodal Model
@@
With the introduction of GPT-4V (OpenAI, 2023), many versatile large multimodal models have been proposed. Commonly, these models aim to handle both text and image tasks simultaneously. They employ a tokenizer and a visual encoder (such as CLIP (Radford et al., 2021)) to encode textual and visual information separately, then concatenate the encoded vectors into a unified input that is fed into the large language model. Notable LLMs include LLaVa (Liu et al., 2024b), Qwen-VL-Max (Bai et al., 2023b), Gemini (Team et al., 2024), InternLMX Composer-VL (Zhang et al., 2024a), and GPT-4o. These models have greatly succeeded in general multimodal tasks, such as OCR, visual question answering, image captioning, multimodal reasoning, etc. Our CMM-Math aims to provide a detailed and comprehensive evaluation of the mathematical multimodal reasoning capabilities of these models. Unlike training methods for general-purpose large multimodal models, we propose a three-stage training method specifically for the mathematical domain.
Table 1. Key statistics of CMM-Math.
@@
Table 2. Comparison with existing datasets. EN and CN mean English and Chinese.
Multimodal Training Evaluation Language #Number Type
MATH (Hendrycks et al., 2021a) EN 12,500 Fill-in-the-blank questions
GSM8K (Cobbe et al., 2021) EN 8,500 Fill-in-the-blank questions
MathQA (Amini et al., 2019) EN 37,259 Fill-in-the-blank questions
MMMU (Yue et al., 2024) EN 11,500 Choice and fill-in-the-blank questions
MATHVISTA (Lu et al., 2024) EN 6,141 Choice and fill-in-the-blank questions
We-Math (Qiao et al., 2024) EN 6,500 Choice questions
MATH-V (Wang et al., 2024b) EN 3,040 Choice and fill-in-the-blank questions
CMM-Math (Ours) CN 28,249 Choice, fill-in-the-blank, yes-no and analysis questions
3. Dataset
3.1. Dataset Construction
Due to the lack of Chinese multimodal mathematical datasets, we collect and release the CMM-Math dataset, constructed with three main steps: data collection, data cleaning, and data annotation.
During the data collection phase, we gathered over 10,000 real exam papers covering 12 grades of mathematics test problems, from primary school to high school in China. Each exam paper includes various types of questions such as multiple-choice, fill-in-the-blank, and analysis problems. These questions contain both visual and textual information, along with their answers, solutions, grades, and question types. Since the exam papers are in PDF format, we used the Mathpix API [^1] https://mathpix.com to extract the text and images into markdown format and downloaded the extracted images locally.
During the data cleaning phase, we first convert the problems from the markdown text into a JSON format, including fields such as question type, modality, images, question, options, answer, and solution. Then, we check and correct issues related to text and image recognition, such as text or mathematical formulas that are incorrectly identified as images, and images that are misrecognized. Finally, we detect and resolve potential issues within each field, including mismatched question types and overly long parsing.
During the data annotation phase, we ask students to check the problems with the format and context to ensure their quality. Then, we divide the problem into 13 different subjects. Due to the large amount of data and the high cost of manual annotation, we chose three large models, GPT-4o, Gemini, and Qwen-VL-max, to vote on the topics and assign them to the topics based on the principle of minority obeying majority.
Table 3. Detailed statistics of CMM-Math datasets.
@@ 3.1. Dataset Construction
3.2. Dataset Analysis
Our CMM-Math dataset contains 28,069 problems with rich textual and visual information (See Table 1 and Table 3). It contains 21,200 textual problems and 6,869 multimodal problems, divided into choice, fill-in-the-blank, yes-no, and analysis problems. We also split the data into 12 levels, corresponding to the basic education stage from the first grade of primary school to the third grade of high school, to ensure the dataset’s applicability and reference value in educational practice. CMM-Math contains 13 knowledge points, covering most of the mathematical fields encountered in middle and high school, particularly logic, algebra, counting, arithmetic, combinatorics, graph theory, topology, statistics, solid geometry, metric geometry, analytic geometry, descriptive geometry, combinatorial geometry, and transformation geometry.
@@ 3.2. Dataset Analysis
3.3. Comparison with Existing Datasets
@@
To clarify the characteristics of our CMM-Math, we compare it with multiple existing mathematical benchmarks, including textual datasets (e.g., MATH, GSM8K, MathQA) and multimodal datasets (e.g., MMMU, MATHVISTA, MATH-V, and We-Math), as shown in Table 2. First, most of these datasets are in English while we focus on Chinese to measure LMMs more comprehensively. Second, our CMM-Math is a multimodal dataset for mathematical reasoning, where each problem may contain multiple images, designed for LMMs. Textual MATH and GSM8K datasets are used to evaluate the performance of LLMs. Third, our dataset contains the evaluation and training datasets. The existing multimodal datasets evaluate the power of LMMs while ignoring how to improve their performance. Fourth, our datasets are diverse, with rich subjects and problem types across 12 grades, which can more deeply distinguish models’ performance in different mathematical contexts. CMM-Math provides 11 knowledge points and multiple problem types, including choice, yes-no, fill-in-the-blank, and analysis. In addition, the CMM-Math problems are derived from real math tests in primary, middle, and high school, covering a wider range of grades than the GSM8K, which only includes primary school knowledge. CMM-Math can better test the problem-solving ability of the model at different stages. Finally, our dataset has detailed analysis and rich question stem content, which can enhance the mathematical reasoning ability of the model.
Figure 3. The diagram of our Math-LMM framework.
4. Our Proposed Method